# Multilingual Performance of a Multimodal Artificial Intelligence System on Multisubject Physics Concept Inventories

### v1

@@ IV.3 Performance by subject area and image usage
The performance on questions with redundant images is similar to or better than text-only questions for all subject areas, except for LAB and RELA inventories. However, in all subject areas, the AI performs worse on questions that require image interpretation than on text-only questions. RELA, OPT, MECH, MATH, and AST exhibit especially large gaps. Furthermore, while QUANT as a whole was not among the worst performing subject areas on image-based questions, QMVI, which is entirely image-based and consists predominantly of required-image problem types, was the worst-performing of all surveys in English. In line with previous research on the topic, this suggests that visual interpretation is one of the major weaknesses of the studied AI system.
| Required | Text | Unneeded | ||
| image | only | image | ||
| Subject | %Correct | %Correct | %Correct | %Correct |
| AST | 81 | 55 | 91 | 100 |
| EM-C | 53 | 52 | 61 | 100 |
| EM-F | 57 | 52 | 73 | 69 |
| LAB | 37 | 27 | 45 | 17 |
| MATH | 69 | 52 | 90 | - |
| MECH | 59 | 44 | 84 | 80 |
| OPT | 53 | 29 | 88 | 81 |
| QUANT | 64 | 55 | 66 | 79 |
| REAS | 75 | 54 | 76 | 87 |
| RELA | 75 | 0 | 85 | 67 |
| THERM | 77 | 72 | 79 | 77 |
| Overall | 63 | 48 | 81 | 79 |
IV.4 Scores by question and language
| Answers correct in EN | ||||
| 0/3 | 1/3 | 2/3 | 3/3 | |
| Language | 19% (284) | 9% (131) | 9% (128) | 64% (955) |
| AR | 17% (16) | 8% (4) | 50% (6) | 85% (50) |
| BN | 5% (7) | 0% (2) | 11% (3) | 61% (18) |
| CA | 14% (7) | 0% (2) | 78% (3) | 93% (18) |
| CS | 17% (8) | 8% (4) | 76% (7) | 88% (33) |
| DE | 14% (26) | 36% (15) | 41% (9) | 87% (46) |
| EL | 14% (36) | 43% (17) | 52% (9) | 83% (49) |
| ES | 18% (102) | 32% (42) | 56% (45) | 88% (227) |
| FA | 20% (18) | 33% (6) | 28% (6) | 59% (26) |
| FI | 24% (33) | 48% (22) | 48% (11) | 90% (85) |
| FIL | 10% (7) | 17% (2) | 44% (3) | 85% (18) |
| FR | 23% (42) | 42% (12) | 43% (21) | 89% (86) |
| HE | 10% (7) | 17% (2) | 67% (3) | 67% (18) |
| HI | 14% (7) | 0% (2) | 11% (3) | 74% (18) |
| HR | 3% (11) | 33% (6) | 67% (4) | 92% (42) |
| HU | 19% (7) | 17% (2) | 56% (3) | 93% (18) |
| ID | 25% (29) | 48% (16) | 44% (15) | 84% (107) |
| IS | 14% (7) | 0% (2) | 67% (3) | 93% (18) |
| IT | 17% (18) | 22% (6) | 56% (6) | 92% (26) |
| JA | 16% (47) | 31% (14) | 35% (17) | 83% (163) |
| MS | 9% (29) | 37% (10) | 52% (9) | 82% (40) |
| NL | 14% (7) | 0% (2) | 56% (3) | 89% (18) |
| NO | 10% (7) | 33% (2) | 56% (3) | 98% (18) |
| PA | 10% (7) | 0% (2) | 56% (3) | 20% (18) |
| PL | 29% (7) | 17% (2) | 67% (3) | 98% (18) |
| PT | 26% (44) | 37% (20) | 48% (14) | 84% (125) |
| RU | 5% (7) | 17% (2) | 67% (3) | 78% (18) |
| SK | 14% (7) | 17% (2) | 67% (3) | 85% (18) |
| SR | 67% (1) | 33% (3) | 0% (1) | 70% (19) |
| SV | 18% (49) | 36% (29) | 44% (18) | 84% (122) |
| TA | 5% (7) | 50% (2) | 22% (3) | 26% (18) |
| TH | 17% (21) | 25% (8) | 24% (7) | 59% (76) |
| TR | 15% (41) | 13% (13) | 45% (11) | 84% (31) |
| UK | 17% (10) | 0% (2) | 33% (1) | 74% (13) |
| ZH | 21% (46) | 43% (24) | 40% (15) | 73% (153) |
| Overall | ||||
| non-EN | 18% (725) | 33% (301) | 48% (274) | 82% (1771) |
| Overall | ||||
| non-EN | ||||
| Text only | 16% (51) | 52% (18) | 65% (51) | 84% (759) |
Table 10 can be used to assess if the AI struggles with the same questions across different languages. One indicator for “struggle” is if the answers for the three independent attempts agree on the correct choice: how consistent is the AI system in its correct answer choices? The by far largest number of inventories was available in English (all except for the TUG-K2.6) and we referenced the performance to questions given in English. There were 955 individual questions across the inventories for which the AI response was correct for all three responses in English. For 128 questions the AI had two out of the three responses correct. There were 131 questions with one of three correct responses and 284 with all three answers incorrect.
Table 10 shows that questions that are difficult in English tend to remain difficult in the other languages and vice versa (with the earlier mentioned exception of Punjabi and Tamil). For example, among the 102 questions for which all three English responses were incorrect, only 18% of responses in the second most common language, Spanish (ES), were correct. This increases to 32% for the 42 questions with 1/3 correct in English, 56% for the 45 questions with 2/3 correct in English, and 88% for the 227 questions where all three English responses were correct. Similar trends can be seen for the other languages and the percentages for all non-English language questions went from 18% (0/3 in EN), to 33% (1/3 in EN), to 48% (2/3 in EN), and 82% (3/3 in EN).
As discussed earlier, part of this could have been due to lower performances for questions that contain images that pose a challenge for the AI invariably of the language. However, the same trend holds if you look at the “text only” questions. While the overall performance increased (see Table 9), the correlation between questions being answered correctly in English and other languages remains. Correct responses for non-English questions went from 16% (0/3 in EN), to 52% (1/3 in EN), to 65% (2/3 in EN), and 84% (3/3 in EN).
| Incorrect | different | |||
| answers | or same | All | EN | Non-EN |
| 3 | 3 different | 8% (98) | 6% (18) | 9% (80) |
| 3 | 3 the same | 53% (617) | 61% (173) | 50% (444) |
| 3 | 2 the same | 38% (448) | 33% (93) | 40% (355) |
| 2 | 2 different | 34% (197) | 28% (37) | 36% (160) |
| 2 | 2 the same | 66% (376) | 72% (94) | 64% (282) |
Well-designed assessment instruments for teaching have multiple-choice distractors that probe particular student misconceptions. Given that we drew our questions from the PhysPort library with many instruments having been designed carefully, we would expect that for our data set students would gravitate to particular incorrect answers. As such, it is interesting to consider if the AI was gravitating toward specific incorrect answers. Table 11 shows the questions that had two or three incorrect answers for the set of the three AI responses. It then shows if the selected multiple-choice items were the same or different. The percentages for the full data set as well as the data set split into English and non-English responses show that the AI frequently picked the same incorrect multiple-choice items. While not all of the assessments analyzed in this study had five possible choices, the majority did, and we can use this as a reference. If the probability of picking a particular incorrect answer was random we would get a 75% probability of having different answers for two incorrect answers and a 25% chance of having the same answer. For our data this was flipped, 66% of all, 72% of English, and 64% of non-English questions had the same two incorrect choices. The same can be seen for cases where all three answers were incorrect. If the likelihood of a particular incorrect response was random, in 6.25% of cases all three would be the same, in 37.5% of cases all three would be different, and in 56.25% of cases two responses would be the same. As shown in Table 11, in reality, the AI responses are far from random and gravitate toward specific incorrect answers (e.g. 53%, corresponding to 618 questions, instead of 6.25% had all three incorrect responses being the same).
IV.5 Scores by language and language switching
Out of 36 tests that were translated, English was the best performing language in 27 (75%) of cases. Table 12 shows the relative performance by language, compared to English. Overall performance does depend on the language of the questions (nominal language). It is notable that for Catalan, Finnish, French, Italian, Norwegian, and Polish, the performance is comparable, with a number of languages close behind. On the other hand, for nominal languages such as Farsi, Hebrew, Hindi, and most notably Punjabi, the AI exhibits much lower performance.
| AR | BN | CA | CS | DE | EL | ES | FA | FI | FIL | FR | HE | HI | HR | HU | ID | IS | IT | JA | MS | NL | NO | PA | PL | PT | RU | SK | SR | SV | TA | TH | TR | UK | ZH | |
| ADT | - | - | - | - | - | - | 1.04 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 1.06 | - | - | - | - | - |
| BEMA | - | - | - | - | - | - | 0.75 | - | - | - | - | - | - | - | - | - | - | - | 0.80 | - | - | - | - | - | 0.87 | - | - | - | 0.77 | - | - | - | - | 0.67 |
| CCI | - | - | - | 0.95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CDPA | - | - | - | - | - | - | 1.20 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CSEM | - | - | - | - | - | - | 0.98 | - | - | - | - | - | - | - | - | 0.69 | - | - | - | 0.98 | - | - | - | - | - | - | - | - | 0.83 | - | - | - | - | - |
| CTSR | - | - | - | - | - | - | 0.85 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.73 | 0.89 | - | - | - | - | - |
| DIRECT | - | - | - | - | 0.91 | 1.07 | 0.67 | - | 1.05 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 1.00 | - | - | - | - | 1.19 |
| DS | - | - | - | - | 0.81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| ECA | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.91 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCA | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 1.00 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCS | - | - | - | - | - | - | - | - | 0.95 | - | - | - | - | - | - | 0.95 | - | - | - | - | - | - | - | - | - | - | - | - | 0.90 | - | - | - | - | - |
| FCI | 0.84 | 0.56 | 0.97 | 0.90 | 0.98 | 0.95 | 0.98 | 0.76 | 0.97 | 0.85 | 1.02 | 0.73 | 0.71 | 0.97 | 0.97 | - | 0.95 | 1.05 | 0.87 | 0.82 | 0.90 | 1.00 | 0.29 | 1.06 | 1.08 | 0.81 | 0.90 | - | 0.94 | 0.32 | 0.68 | 0.95 | - | 0.82 |
| FMCE | - | - | - | - | - | - | 1.13 | - | - | - | - | - | - | - | - | 1.04 | - | - | 0.54 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| FTGOT | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.87 | - | - |
| GECI | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 1.00 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| HTCE | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.89 |
| IBCDC | - | - | - | - | - | - | - | - | - | - | 1.03 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| IBCM | - | - | - | - | - | - | - | - | - | - | 0.89 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LPCI | - | - | - | - | - | - | 0.95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MBT | - | - | - | - | 0.94 | 1.00 | 0.91 | 0.62 | 1.21 | - | 0.94 | - | - | - | - | - | - | 0.88 | 1.00 | 0.88 | - | - | - | - | 0.79 | - | - | - | - | - | - | 0.85 | - | - |
| MWCS1 | - | - | - | - | - | - | 0.96 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.44 | - | - | - |
| MWCS2 | - | - | - | - | - | - | 0.91 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| NGCI | 0.87 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMCA5 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.90 | - | - | - | - | - | - | - | - | - |
| QMCS | - | - | - | - | - | - | - | - | 0.94 | - | - | - | - | - | - | - | - | - | 0.90 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMFPS | - | - | - | - | - | - | 0.82 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QPCS | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.57 | - | - | - |
| RFCI | - | - | - | - | - | - | - | - | 1.00 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RKI | - | - | - | - | - | - | - | - | - | - | 0.95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| SPCI | - | - | - | - | - | - | 0.98 | - | - | - | - | - | - | - | - | - | - | - | 0.90 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| STPFaSLSh | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.81 | - | - | - | - | - | - | - | - | 0.84 | - | - | - | - | - | - | - | - | 0.68 |
| TCE | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.96 | - | - | - | - | - | 0.88 | - | - | - | - | - | - | - | - | 0.67 |
| TCS | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.75 | - | - | 0.87 |
| TOAST | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.97 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TUG-K | 0.86 | - | - | - | 1.09 | 0.79 | 1.09 | - | 1.09 | - | 1.21 | 0.78 | - | - | - | - | - | - | - | - | - | - | - | - | 1.19 | - | - | - | 0.91 | - | - | - | 0.81 | 0.81 |
| TUV | 0.97 | - | - | - | - | - | 1.03 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| Average | 0.89 | 0.56 | 0.97 | 0.92 | 0.95 | 0.95 | 0.95 | 0.69 | 1.03 | 0.85 | 1.01 | 0.75 | 0.71 | 0.94 | 0.97 | 0.90 | 0.95 | 0.97 | 0.88 | 0.90 | 0.90 | 1.00 | 0.29 | 1.06 | 0.94 | 0.81 | 0.90 | 0.73 | 0.91 | 0.32 | 0.61 | 0.89 | 0.81 | 0.83 |
It is unclear in which language the AI system internally processes the problems, however, the language of the output may be a good indicator, which exhibits code-switching. As seen in Table 14, for some languages, the AI performs better when the output is either entirely in English or mixed with English. Languages like Bengali, Persian, and Punjabi have a marked increase in accuracy when the AI switches to English, implying difficulties in processing the original language of the inventories. Meanwhile, languages like English, Spanish, and French demonstrate more balanced performance, with nominal language usage still resulting in high correctness scores. If the performance of the AI system for a particular question in English is used as a measure of difficulty, Table 13 indicates that the AI stays in the nominal language more often for easier questions. For all non-English questions combined, it does so for 24% of the questions that are deemed easier in English (3/3 correct answers in English). While not a strong signal, it is markedly more than 15-17% for the questions where at least one response in English was incorrect.
| Answers correct in EN | ||||
| Language switch | 0/3 | 1/3 | 2/3 | 3/3 |
| Switch to English | 69% | 69% | 69% | 63% |
| Mixed Language | 14% | 17% | 15% | 12% |
| Nominal Language | 17% | 15% | 17% | 24% |
Clustering emerges if we consider how often different languages remain in their nominal language versus switching to English or mixing languages; Fig. 7 shows a heat map of the cosine-similarity between the tuples (%nominal language, %switching to English, %mixed language) for each language. Spanish and Portuguese form a cluster that in the majority of cases stays within the nominal languages; the two languages are also part of a larger cluster of commonly spoken European languages, namely German, Italian, Dutch, and French. The next cluster includes several Scandinavian and Eastern European languages. Finally, a large cluster of mostly Asian and Arabic languages emerges, which with the notable exceptions of Polish and Icelandic uses non-Latin script. Thus, from a linguistic standpoint, we can group languages based on their dependence on language switching: those that remain robust in their native form, those that need English to boost correctness, and those that benefit from mixing strategies. This language-based clustering mirrors the subject-area clustering by revealing underlying differences in how the AI system processes and understands content under different linguistic and visual conditions.
| Nominal language | Switch to English | Mixed languages | |||||
| Language | %Correct | % | %Correct | % | %Correct | % | %Correct |
| AR | 60 | 6 | 89 | 92 | 58 | 2 | 83 |
| BN | 39 | 2 | 100 | 98 | 38 | 0 | |
| CA | 67 | 19 | 94 | 60 | 57 | 21 | 68 |
| CS | 69 | 3 | 100 | 58 | 54 | 40 | 89 |
| DE | 56 | 19 | 70 | 40 | 49 | 41 | 56 |
| EL | 52 | 8 | 63 | 80 | 50 | 11 | 61 |
| EN | 72 | 100 | 72 | - | - | - | - |
| ES | 62 | 59 | 63 | 33 | 62 | 8 | 52 |
| FA | 40 | 1 | 0 | 98 | 40 | 1 | 100 |
| FI | 66 | 4 | 86 | 60 | 65 | 36 | 64 |
| FIL | 59 | 0 | 100 | 59 | 0 | ||
| FR | 63 | 37 | 74 | 43 | 55 | 20 | 60 |
| HE | 47 | 1 | 100 | 97 | 46 | 2 | 67 |
| HI | 49 | 9 | 75 | 90 | 47 | 1 | 0 |
| HR | 69 | 11 | 81 | 64 | 69 | 25 | 66 |
| HU | 67 | 9 | 100 | 70 | 65 | 21 | 58 |
| ID | 67 | 26 | 71 | 59 | 67 | 15 | 60 |
| IS | 66 | 0 | 100 | 66 | 0 | ||
| IT | 57 | 37 | 71 | 53 | 48 | 10 | 47 |
| JA | 63 | 8 | 87 | 83 | 60 | 8 | 72 |
| MS | 50 | 0 | 100 | 50 | 0 | ||
| NL | 62 | 34 | 68 | 39 | 46 | 27 | 79 |
| NO | 69 | 13 | 75 | 71 | 62 | 16 | 93 |
| PA | 20 | 2 | 100 | 93 | 19 | 4 | 0 |
| PL | 73 | 0 | 100 | 73 | 0 | ||
| PT | 64 | 56 | 73 | 34 | 50 | 10 | 63 |
| RU | 56 | 8 | 57 | 84 | 55 | 8 | 57 |
| SK | 62 | 0 | 76 | 62 | 24 | 64 | |
| SR | 62 | 0 | 100 | 62 | 0 | ||
| SV | 60 | 13 | 65 | 56 | 59 | 31 | 58 |
| TA | 22 | 1 | 100 | 87 | 19 | 12 | 36 |
| TH | 47 | 7 | 67 | 86 | 44 | 7 | 55 |
| TR | 40 | 18 | 51 | 64 | 39 | 18 | 35 |
| UK | 45 | 3 | 0 | 79 | 45 | 18 | 50 |
| ZH | 58 | 2 | 57 | 97 | 58 | 1 | 20 |
| Overall | 63 | 46 | 71 | 45 | 55 | 9 | 60 |
V Discussion
The results presented here highlight several noteworthy patterns in the AI’s performance across different physics inventories, languages, and question formats. Unlike many previous studies where inventories were provided as text-only materials, this study supplied the AI with a screenshot of the question the way a student would see it. This richer, more realistic presentation included the actual layout, images, and language-specific formatting, effectively increasing the complexity of the input. The AI’s decision-making process—whether to remain in the nominal language, switch entirely to English, or produce a mixed-language response—was an emergent behavior rather than an externally imposed constraint. The prompt itself was given in English, but the question content was often in another language, with images embedded in the screenshot. Thus, the AI’s language choices and solution correctness provide insights into its internal reasoning and capability to handle diverse formats.
Across the majority of inventories and subject areas (with laboratory skills being the exception), the AI achieves higher scores and performs better than the average undergraduate student post-instruction, based on published results. It needs to be emphasized that the instruments are research-based and thus carefully designed with regard to psychometric properties, which is reflected in their score distributions:
The test items aim for medium difficulty, resulting in average scores around 50%.
The test items are designed for high discrimination, resulting in broad score distributions.
Thus, the finding that the AI performs better than the average published results, in broad strokes, means that it performs better than 50% on these instruments. It also means that there will almost always be students who outperform the AI. Whenever a study included results for graduate-student or post-doctoral populations, these clearly outperformed the AI.
Looking at the presence and function of visual representations in questions, we can see a clear pattern. The AI performed better on text-only questions and questions with redundant visual representations, compared to questions where visual interpretation of images was necessary for solving the task correctly. This was true for each individual subject area, as well as across the entire dataset. The average difference in performance between required image and text-only questions was 33 percentage points, with some subject areas exhibiting much higher gaps (see Table 9). This corroborates previous findings on the topic, which suggest that the AI’s vision abilities present a major weakness and hamper its ability to engage with physics tasks that require the interpretation in graphical or picture formats commonly found in physics [24, 31]. For at least the present time, instructors who want to discourage students from using AI to do their homework, including required information in some kind of graphical or picture (or even video) format is good advice.
On the language side, the results show a complex interplay between linguistic capability and the AI’s self-directed switching behavior. With the exception of Spanish and Portuguese, the AI switches away from the nominal language in the majority of cases. This outcome likely points to limitations in the AI’s mastery of certain languages, where staying in the original language of the prompt results in lower accuracy. Conversely, in languages such as German, French, Spanish, and Portuguese, the AI performs better on questions where it stays in the nominal language, which may have at least two possible explanations:
The AI is more adept at these languages or that their training data better captures the relevant scientific vocabulary and discourse, or
The AI switches languages for the problems where it already “struggles,” and then performs worse on those than on the ones where straightforward inferences can be found in the nominal language.
The fact that the AI autonomously decided to switch or mix languages further suggests that it was “aware,” at least statistically, of its varying strengths and weaknesses across linguistic contexts.
Our results show that questions that are difficult in English are likely difficult in other languages as well. Either the AI switched to English in its reasoning or, when staying in the nominal language, as frequently seen in Spanish, similar questions were difficult (or easy) for the AI even though the training data sets in the different languages were not the same. Like students, the AI gravitated toward particular incorrect solutions. How the level of difficulty for a specific question type is different for students versus AI is intriguing. When going astray, do students and AI gravitate toward the same incorrect answers? Are the preferred incorrect AI selections a reflection of issues in the training data or just more challenging for the logic built into GPT-4o? While this study does not answer these questions, it provides a stepping stone for future research. Our study indicates we can identify questions that are consistently challenging for AI across languages. Designers and instructors could specifically look at assessment designs that are challenging for AIs. The most obvious is the inclusion of images but this work clearly shows text-based questions can also pose challenges across languages for a multimodal AI system.
Overall, the study’s methodology – providing authentic screenshots rather than merely text and having the AI itself choose its linguistic mode – brings us closer to understanding how such systems would behave in real-world educational settings. We see that certain subject matters are intrinsically more challenging, that language understanding varies widely, and that the AI’s own adaptive language strategies can partially compensate for these differences. Nonetheless, large gaps and outliers remain, pointing to areas where improvements in training data, prompt design, or model architecture are needed. Future work should focus on exploring how to guide the system’s language-switching decisions more effectively, refining the visual understanding components, and ensuring more balanced coverage across different languages and conceptual domains.
VI Limitations
[…]

### v3

@@ IV.4 How does the presence of images influence the performance of the AI system?
When examining individual subject categories, GPT-4o consistently performs worse on items that require image interpretation than on text-only items. Relativity, Optics, Mechanics, and Mathematics (which had no unneeded image items), as well as Astronomy, exhibited especially large performance gaps. Furthermore, while QP as a whole was not among the lowest-performing subject categories on image-based problem types, QMVI — entirely image-based and composed predominantly of required-image items — was the second worst-performing inventory in English (32%). In line with previous research on the topic [36], these findings suggest that visual interpretation is one of GPT-4o’s major weaknesses.
V Discussion and future research directions
Unlike many previous studies where inventories were provided as text-only materials, this study provided the AI with a screenshot of the item as it would appear to a student. Using authentic screenshots rather than isolated text offers a closer approximation of the model’s behavior in realistic educational settings, especially in multilingual and multimodal contexts. However, this approach also increases the complexity of the input and the analytical demand for interpreting the model outputs.
Across the majority of inventories and subject categories (with Laboratory skill being the exception), GPT-4o achieves higher scores and outperforms the average reported student post-instruction undergraduate student. It needs to be emphasized that the instruments are research-based and thus carefully designed with regard to psychometric properties, which is reflected in their student score distributions: (i) The inventory items aim for medium difficulty, resulting in average student scores around 50%, and (ii) The items are designed for high discrimination, leading to broad score distributions. Thus, the finding that the AI performs better than the average student on these inventories broadly means that it performs above 50% on these instruments. It also means that in almost all cases, some students in the dataset still outperformed GPT-4o. Furthermore, in studies where results for graduate or post-doctoral populations were available, these outperformed GPT-4o.
It is important to note that even when GPT-4o outperformed the average student, this does not mean that it necessarily exhibited a strength and difficulty profile similar to that of a well-performing student. Our findings support previous research (e.g., [37]) and suggest that image interpretation is one common GPT-4o difficulty that is not typically seen in students.
This highlights the importance of carefully considering the role of AI in physics assessment: while it can match or exceed undergraduate averages on concept inventories, its performance should not be mistaken for human expert-like reasoning or deep conceptual understanding. For example, a test on which AI performs poorly might not necessarily be difficult for students, and vice versa. Much depends on the type of tasks involved and how they align with the AI’s and students’ strength and difficulty profiles.
If AI systems like GPT-4o will play a role in the pipeline of assessment design and validation, the role will likely be different from that of a student or a domain expert. While there exists some research on making an LLM behave like a student with a certain difficulty profile [157], more work is needed to establish the feasibility of such use at scale and across different subject areas. On the other hand, because of the increasingly important role AI systems will likely play in physics education, it is worthwhile considering developing assessments specifically for those AI systems that are going to be used in educational contexts. Such assessment would provide curriculum developers and instructors with even more pertinent information about these systems’ capabilities and potential in educational contexts. Exactly what such assessment would look like remains unclear. However, it is reasonable to expect that it will be developed in response to the practical needs and demands of the new educational landscape.
On the linguistic side, our results reveal a complex relationship between performance and language. The overall pattern that emerged is that GPT-4o performed better in most cases when tasks were given in English. While some other languages — mostly European and Latin-script — exhibited roughly similar performance, others show significantly lower results. One contributing factor for this difference could be that GPT-4o had to handle English language prompts with screenshots using other language scripts, a task that is presumably harder than just dealing with one language or script. Understanding this better would need to be explored further in a future study. This is particularly importantnt as it has implications for accessibility and utility of AI tools for non-Western language-speaking populations and may risk perpetuating or even exacerbating disparities in access to educational and technological resources worldwide.
Our results also indicate that tasks found to be challenging in English tend to remain so when posed in other languages. Notably, the model often produces consistent patterns of incorrect responses, even though its training data likely differs across languages. This observation raises questions about whether these errors stem from biases in the training data or reflect deeper limitations in the model’s inference process. While our study does not resolve these issues, it highlights that certain items consistently challenge the model across languages. These findings may inform future research on assessment designs that account for the limitations of multimodal AI systems, whether the items are visual or text-based.
Examining the presence and function of visual representations in inventory items reveals a clear pattern. The AI performed better on text-only items and items with redundant visual representations, compared to items where visual interpretation of images was necessary for solving the task correctly. This corroborates previous findings on the topic, suggesting that the AI’s visual reasoning remains a major weakness, hampering its ability to engage effectively with graphical or picture formats common in physics tasks [31, 37]. Anecdotal evidence in recent months suggests that, compared to GPT-4o, GPT-o3 and o4-mini [158] perform better on physics tasks, including numerical ones, but often still struggle with the interpretation of figures.
There may be other variables, beyond subject area, language and image presence, which influence the performance of GPT-4o and other AI systems on physics tasks. The mathematical complexity of a task and the presence of redundant information are just two of the many possible aspects that could influence AI’s ability to solve a task. More research is needed to develop a better sense of what makes a physics task difficult for an AI system to solve.
Based on our findings, we suggest that future efforts to improve AI performance on physics tasks should focus on enhancing the multimodal processing components and achieving more balanced performance across languages and conceptual domains.
VI Implications for instruction
Our study is decidedly exploratory and does not aim to provide direct advice to instructors on how to implement GPT-4o in education. Still, there are some broad findings that emerged from our exploration of GPT-4o’s performance on physics concept inventories, which can be informative for physics instructors and curricular developers.
We expect that in the future, multimodal AI systems like GPT-4o will continue to influence the space of physics education through at least two mechanisms. First, their accessibility and availability and their relatively high capabilities make them attractive for learners. They will likely continue to be used by students to help them with physics tasks. Second, AI tools are likely to see increased uptake by educational institutions to support students in their learning as well as to support instructors in their teaching and administrative tasks.
It is important to note once more that while GPT-4o, on the surface, exhibits performance that is numerically better than university students’ post-instruction average, a closer look reveals important caveats with this simple interpretation. If students want to use GPT-4o productively and responsibly, they should be aware of its limitations. We believe that instructors should inform students of these drawbacks to mitigate the risks of over-reliance on AI tools, and foster a critical perspective on the outputs these tools generate. This should arguably become one of the newly emerging instructional goals because such skills will remain useful even as students leave education and enter the workforce. As AI becomes part of our everyday and work, evaluating its output becomes an important skill, which cannot be learned by always outsourcing physics reasoning to an AI. Exposing AI’s drawbacks to students can thus also serve as a motivation for students to engage more deeply in learning physics. Our research can help inform such efforts. For example, we have shown that GPT-4o’s performance is not equally good across all subject areas of physics and that it often struggles when prompted in non-Western languages. Furthermore, a major drawback is its ability to interpret images.
However, because of the incredible pace of AI progress, sooner or later, physics educators will have to contend with the question of what remains meaningful to teach, when AI systems perform well on many tasks that were previously squarely in the domain of human physics experts. The physics education community will likely need to seriously reflect on whether physics curricula, which have in many cases remained nearly unchanged for decades, should evolve to better reflect the new reality. If we conclude that the kind of conceptual understanding that is being tested by research-based concept inventories is still valuable and important for our students, then these assessments will likely continue to play an important role in evaluating whether our students have reached the desired learning objectives. In such a case, they should arguably be administered so that students cannot use AI to help them.
In making these important decisions, the physics education community will likely also need to address the following (and other) questions on a continuous basis: What are the foundational skills that we should not routinely outsource to AI? What are the central tasks that students first need to master themselves, so that they can later judiciously and responsibly outsource them to AI? How do we ensure that access to top-performing AI does not generate or worsen existing divides based on students’ economic or ethnic background?
VII Limitations
This study is decidedly exploratory and empirical. The preparation of item images was done manually, and some manual cleanup of the data was required. Given more than 3,600 images and the random oddities occurring in over 14,000 solutions generated by a probabilistic system, clerical errors cannot be excluded. Additionally, each screenshot of inventory items was iterated only three times: given the stochastic nature of LLM outputs, this introduces variability that limits the results’ generalizability. The study also did not consider the quality of the concept inventories’ translations, and lower scores may be due to incorrect or confusing translations. AI’s performance might be dependent on prompts, and the ones used for our study (shown in Figs. 8 and 9 of Appendix A) may not be the best choices. Future studies may look into whether and how prompt-engineering techniques might improve (or worsen) AI’s performance.
Furthermore, it is unclear to what extent the use of an English prompt, combined with inventory text in the nominal language of each inventory, influenced performance across languages and contributed to language-switching behavior; Appendix B discusses some of the preliminary observations. Future research may explore this further by varying the prompting approach, for example, by prompting the system entirely in the nominal language. However, due to the unreliability of structured outputs with non-English prompts (see Appendix A), such approach would likely require alternative output formats, more human involvement in the coding of the answers, as well as accurate translations of the prompts into multiple languages.
The model’s visual encoder transforms the pixel data into a latent representation that encapsulates both textual and graphical information. This has the side effect that there is no standalone raw OCR (optical character recognition) output that we could compare to the original text in the various Latin and non-Latin scripts used in the study. One possible explanation for the difference in performance between languages could be incorrect recognition of non-Latin characters.
Furthermore, it should be noted that we did not score the correctness of the physics reasoning in the AI-written explanations. Future studies could explore this in more detail, potentially using the data collected in this study [159] (as an example, see Fig. 11 in Appendix A) to evaluate the model’s reasoning separately from its final answers. An example of how this could be done can be found in [36].
A possible stumbling block for the AI can be the processing of graphical information unrelated to a physics concept or language. For example, for item 7 of the FCI, the multiple-choice options are embedded in a graphic showing a ball swinging in a circular path and are not listed separately in the text. Although the AI frequently gave physically reasonable explanations, across 32 languages and 96 total answers, it selected the correct embedded option only once. For no obvious reason, it chose instead the same incorrect answer an eyebrow-raising 88 times. Once again, this finding aligns with previous studies showing that the graphical layout of images and the spatial arrangement of answer options can play an important role in the AI’s selection of answers [37].
The performance of GPT-4o, as measured in our study represents a momentary snapshot of one model’s capabilities in early 2025. It is very likely that future models will outperform GPT-4o. However, to keep track of such developments, more studies similar to ours will be needed.
Finally, human post-instruction scores were gathered on a best-effort basis, which may introduce additional variability into the comparisons. Moreover, most of the human data came from English-speaking students taking the English versions of the inventories.
[…]